Suppose that the average waiting time for a patient at a physician's office is just over 29 minutes. In order to address the issue of long patient wait times, some physicians' offices are using wait-tracking systems to notify patients of expected wait times. Patients can adjust their arrival times based on this information and spend less time in waiting rooms. The following data show wait times (minutes) for a sample of patients at offices that do not have a wait-tracking system and wait times for a sample of patients at offices with a wait-tracking system.

without with

23 31
62 10
15 13
21 17
32 11
45 35
10 10
26 2
16 11
35 16

(a)Considering only offices without a wait-tracking system, what is the z-score for the 10th patient in the sample (wait time = 35 minutes)?

If required, round your intermediate calculations and final answer to two decimal places.

z-score =  

(b)Considering only offices with a wait-tracking system, what is the z-score for the 6th patient in the sample (wait time = 35 minutes)?

If required, round your intermediate calculations and final answer to two decimal places.

z-score =  

How does this z-score compare with the z-score you calculated for part (a)?
Answer #1

Ans:

z score=(x-mean)/std. dev

without with
1 23 31
2 62 10
3 15 13
4 21 17
5 32 11
6 45 35
7 10 10
8 26 2
9 16 11
10 35 16
Mean= 28.5 15.6
std. dev(s)= 15.71 10.07
z= 0.41 1.93

Without:

z=(35-28.5)/15.71=0.41

with:

z=(35-15.6)/10.07=1.93